OpenStudy (anonymous):

help @ganeshie8 @mathstudent55

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

@Eleven17 ?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

whats the question

ganeshie8 (ganeshie8):

27 = 3 * 3 * 3 45 = 3 * 3 * 5

ganeshie8 (ganeshie8):

thats prime factorization of both the given numbers

ganeshie8 (ganeshie8):

LCM = product of greatest powers of primes

OpenStudy (anonymous):

ok i understand so far

ganeshie8 (ganeshie8):

good :) since 3^3 is large, we take it. and since 5 is there oly in the second number we take it. so LCM = 3^3 * 5 = ?

OpenStudy (anonymous):

185?

ganeshie8 (ganeshie8):

each prime factor you need to see in which number it is greatest, and take it

ganeshie8 (ganeshie8):

3^3 * 5 = 27 * 5 = 135

OpenStudy (anonymous):

ooo so thats the answer ?

OpenStudy (anonymous):

idk what i did wrong

ganeshie8 (ganeshie8):

yes ! what did u get ?

OpenStudy (anonymous):

ooo i seee what i did wrong

ganeshie8 (ganeshie8):

oh 185 ? forget it... must be u did a typo..